# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ Abstract
In this paper we consider symplectic 4-manifolds $(M,\omega)$ with $c_{1}(M,\omega)=0$ which admit a Hamiltonian $S^{1}$-action together with an equivariant Maslov condition on orbits of the group action. We call such spaces special Hamiltonian $S^{1}$-spaces. It turns out that there are no compact special Hamiltonian $S^{1}$-spaces. We classify all exact special Hamiltonian $S^{1}$-spaces and show that all of them admit the structure of a Stein surface. Moreover, we show that any special Hamiltonian $S^{1}$-action on a 1-connected Stein surface $\mathcal{W}_{n}$ is eventually symplectically equivariant to the standard one if $n=0,1$, and at least near the $A_{n}$-string of Lagrangian spheres associated to the standard special Hamiltonian $S^{1}$-action if $n\geq 2$.
@@ 1 Introduction and main results
Hamiltonian $S^{1}$-actions on compact symplectic 4-manifolds have been classified by Karshon in [8] where she proved that all these Hamiltonian $S^{1}$-spaces are Kähler. In this article we consider a type of Hamiltonian $S^{1}$-action $G:S^{1}\to Ham(M,\omega)$ on a symplectic 4-manifold $(M,\omega)$ with an extra property called the Maslov condition, which requests that along any nondegenerate orbit $C$, any $G$-invariant Lagrangian subbundle in the symplectic normal bundle of $C$, treated as a loop of Lagrangian planes along $C$, has zero Maslov number. We call such a group action a special Hamiltonian $S^{1}$-action and the triple $(M,\omega,G)$ a special Hamiltonian $S^{1}$-space. See Definition 2.1 in §2 for the precise definition of a special Hamiltonian $S^{1}$-action. To ensure the Maslov condition is permitted we assume that $(M,\omega)$ is connected with the first Chern class $c_{1}(M,\omega)=0$, and that $G$ acts effectively on $M$.
It turns out that the Maslov condition imposes restrictions on the topology of $M$ as well of the $G$-action.
@@
Let $(M,\omega,G)$ be a connected special Hamiltonian $S^{1}$-space and $G$ acts effectively on $M$ with moment map $h$, then
@@
    the fixed point set $Fix(G)$ of the $G$-action is a discrete set of points (Proposition 3.4).
@@
    Moreover, if $(M,\omega=d\lambda)$ is exact then
    1. (a).
      $Fix(G)\subset h^{-1}(0)$ up to addition of a constant to $h$, $h^{-1}(c)$ is connected for $c\in\mathbb{R}$, and
    2. (b).
      the $G$-action is semi-freeI.e., the stabilizer $stab_{p}:=\{g\in G\mid g(p)=p\}$ is $\{1\}$, $\forall p\in M\setminus Fix(G)$.. (Proposition 3.5)
In this paper we classify all connected exact special Hamiltonian $S^{1}$-spaces $(M,\omega=d\lambda,G)$ with finite fixed point set $Fix(G)$. For the 1-connected case we obtain the following classification results.
@@
Let $(M,\omega,G)$ be a connected, 1-connected exact special Hamiltonian $S^{1}$-space, then the $G$-action is semi-free with $n+1$ fixed points for some integer $n\geq 0$. If $n>0$ then $(M,\omega)$ is symplectically the linear plumbing of $n$ cotangent bundles of spheres $S_{i}\subset M$, $i=1,...,n$, where the union $\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of $G$-invariant Lagrangian spheres. If $n=0$ then $(M,\omega)$ is symplectomorphic to the standard symplectic 4-space.
@@
Let $\omega_{n}$ denote the standard symplectic structure on $\mathbb{C}^{3}$ restricted to $\mathcal{W}_{n}$. It is known that for $n>0$ $\mathcal{W}_{n}$ is topologically the plumbing of $n$ copies of cotangent bundles $T^{*}S^{2}$ of type $A_{n}$ [16], and $\mathcal{W}_{0}$ is symplectomorphic to the standard symplectic 4-space.
@@
Let $\tilde{\textsf{G}}$ denote the $S^{1}$-action on $\mathbb{C}^{3}$ defined by
@@
Then $\tilde{\textsf{G}}$ acts on $(\mathcal{W}_{n},\omega_{n})$ as a special Hamiltonian $S^{1}$-action with $n+1$ Fixed points.
@@
A connected, 1-connected, exact special Hamiltonian $S^{1}$-space $(M,\omega=d\lambda,G)$ with $n+1$ fixed points is symplectomorphic to the Stein surface $(\mathcal{W}_{n},\omega_{n})$. In particular, the symplectic topology of $(M,\omega)$ is determined by the cardinality of the fixed point set $Fix(G)$.
Moreover, together with the following theorem by Wu [16], any $A_{n}$-string of Lagrangian spheres in $(M,\omega)$ is Hamiltonian isotopic to any given $A_{n}$-string of Lagrangian spheres associated to the $G$-action as in Theorem 1.4, up to a composition of Lagrangian Dehn twists along the latter:
@@
Let G denote the linear $S^{1}$-action of $\tilde{\textsf{G}}$ restricted to $\mathcal{W}_{n}$. In Section 5 we will construct an explicit G-invariant $A_{n}$-string of Lagrangian spheres $\textsf{S}:=\cup_{j=1}^{n}\textsf{S}_{j}$ for in $(\mathcal{W}_{n})$. Theorem 1.4 and Theorem 1.5 together implies that up to a symplectomorphism we can identify $(M,\omega,G)$ with $(\mathcal{W}_{n},\omega_{n},G)$ and S in $\mathcal{W}_{n}$ is preserved by the $S^{1}$-action of $G$. The classification of 1-connected special Hamiltonian $S^{1}$-space is then reduced to the classification of semi-free special Hamiltonian $S^{1}$-actions on $(\mathcal{W}_{n},\omega_{n})$.
We have the following result concerning the uniqueness of special Hamiltonian $S^{1}$-actions on $\mathcal{W}_{n}$:
Theorem 1.6.
Let $G$ be any semi-free special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{n},\omega_{n})$.
  1. (i).
    The $S^{1}$-action of $G$ is symplectically equivalent to the $S^{1}$-action of G on some open neighborhood of the $A_{n}$-string of Lagrangian spheres associated to G. I.e., there is a $G$-invariant open neighborhood $U\subset\mathcal{W}_{n}$ of S (resp. the fixed point $(0,0,1)\in\mathcal{W}_{0}$ if $n=0$) and a $C^{\infty}$-embedding $\EuFrak{e}_{\textsf{S}}:U\to W_{n}$ on which $G=\EuFrak{e}_{\textsf{S}}\circ\textsf{G}\circ\EuFrak{e}_{\textsf{S}}^{-1}$ and hence $h=\textsf{h}\circ\EuFrak{e}_{\textsf{S}}^{-1}$, where $h$ and h are the moment maps of $G$ and G respectively.
  2. (ii).
    If $n=0$, $G$ is symplectically equivariant to G on any compactly supported $G$-invariant open domain containing the fixed point $(0,0,1)$ of the G-action on $\mathcal{W}_{0}$.
  3. (iii).
    If $n=1$, then $G$ is symplectically equivariant to G on any compactly supported $G$-invariant open domain containing the $A_{1}$-sphere $S\cong S^{2}$ associated to the G-action on $\mathcal{W}_{1}$.
For the case that $(M,\omega=d\lambda)$ is not 1-connected we assume that $Fix(G)$ is not empty (otherwise the $G$-action is free and the Maslov condition is vacuous). Then the reduced space $M_{red}$ is topologically an oriented noncontractible surface of genus $g\geq 0$ with $b\geq 1$ boundary components, $g+b\geq 2$, and $k:=|Fix(G)|=n+1\geq 1$ marked points.
Theorem 1.7.
Let $(M,\omega,G)$ be a connected exact special Hamiltonian $S^{1}$-space with $c_{1}(M,\omega)=0$ and $n+1$ fixed points, $n\geq 0$. Assume also that $M$ is not 1-connected. Then $M$ admits the structure of a Stein surface, which is topologically obtained by attaching $(2g+b-1)$-pairs of 1- and Lagrangian 2-handlebodies to the boundary of a Stein domain $W_{n}\subset\mathcal{W}_{n}$ diffeomorphic to $\mathcal{W}_{n}$. The homology groups of $M$ are
$H_{m}(M,\mathbb{Z})=\begin{cases}0&\text{ for }m>2,\\ \mathbb{Z}^{2g+b-1+n}\cong H_{2}(W_{n},\mathbb{Z})\oplus\mathbb{Z}^{2g+b-1}&\text{ for }m=2,\\ \mathbb{Z}^{2g+b-1}\cong H_{1}(M_{red},\mathbb{Z})&\text{ for }m=1.\end{cases}$
In particular, $H_{2}(M,\mathbb{Z})$ is generated by $n$ Lagrangian spheres and $12g+b-$ 1 monotone Lagrangian tori.
Remark 1.8 (Motivation on Maslov condition and relation with complexity one Hamiltonian space).
A motivation to study 4-dimensional special Hamiltonian $S^{1}$-spaces is to explore Hamiltonian $S^{1}$-spaces which possess monotone Lagrangian surfaces invariant under the given Hamiltonian $S^{1}$-actions. It turns out that the 4-dimensional exact special Hamiltonian $S^{1}$ spaces $(M,\omega=d\lambda,G)$ that we consider here are centered complexity one spaces as defined in [9] (see also [10, 11, 12]) except that we do not demand the moment maps of these spaces to be proper, as was requested in [9]. Ignoring the condition on properness, the Maslov condition that we imposed here provides a concrete subclass of centered complexity one spaces which contain various types of immersed monotone Lagrangian surfaces, including spheres, tori, and Whitney spheres (immersed spheres with one transversal self-intersection point), etc., all contained in the 00-level set of the moment map associated to the special Hamiltonian $S^{1}$ action. See Remark 4.2 after Proposition 4.1. It will be interesting to see how these Lagrangian surfaces, in addition to their topological types, are classified up to Hamiltonian isotopy.
@@
The rest of this paper is organized as follows: In §2 we give the definition of a special Hamiltonian $S^{1}$-action on a symplectic 4-manifold $(M,\omega)$. As an example we show that an $S^{1}$-subgroup of $SU(2)$ which acts linearly on the standard symplectic 4-space $(\mathbb{R}^{4},\omega_{0})$ meets the special Hamiltonian criteria. Proof of Theorem 1.1 is given in §3 where we also establish various symplectic topological properties of special Hamiltonian $S^{1}$-spaces. Theorem 1.2 is proved in §4. Proposition 1.3 and Theorem 1.4 are verified in §5 where the $\tilde{\textsf{G}}$-action is analyzed and an $A_{n}$-string of Lagrangian spheres is given. This is followed by the proof of Theorem 1.6. Theorem 1.7 is proved in §6.

### v4

@@ Abstract
In this paper we consider symplectic 4-manifolds $(M,\omega)$ with $c_{1}(M,\omega)=0$ which admit a Hamiltonian $S^{1}$-action together with a Maslov condition on orbits of the group action. We call such spaces special Hamiltonian $S^{1}$-spaces (SHam1-spaces) and denote them as $(M,\omega,\Phi,h)$ or $(M,\omega,\Phi)$, where $\Phi:S^{1}=\mathbb{R}/2\pi\mathbb{Z}\to Ham(M,\omega)$, $\Phi(t)=\phi_{t}$, $\phi_{0}=id_{M}$, is a group homomorphism, $h$ is the associated moment map. It turns out that there are no compact SHam1-spaces. To minimize topological complexity we assume that $(M,\omega=d\alpha)$ is an exact, connected open tame manifold with connected level sets $h^{-1}(c)$, $c\in\mathbb{R}$, the gradient vector field $\nabla h$ of the moment map $h$ is complete with respect to some $\omega$-compatible Riemannian metric on $M$, and all orbit spaces $h^{-1}(c)/\Phi$ are homeomorphic to a complete manifold. We also assume that the SHam1-action on $(M,\omega=d\alpha)$ is effective, semi-free, and the fixed point set is finite. SHam1-spaces with the above conditions are called simple. We classify all simple SHam1-spaces and show that all of these spaces admit the structure of a Stein surface. Moreover, we show that, a SHam1-action $\Phi$ on a 1-connected Stein surface $\mathcal{W}_{n}$ is linear near the $A_{n}$-string of Lagrangian spheres, up to a conjugation with a $\Phi$-equivariant diffeomorphism near the $A_{n}$-string of Lagrangian spheres. If $n=0,1$ the then linearity of a SHam1-action $\Phi$ can be further expanded to larger domains by employing suitable Liouville functions.
@@ 1 Introduction and main results
Hamiltonian $S^{1}$-actions on compact symplectic 4-manifolds have been classified by Karshon in [8] where she proved that all these Hamiltonian $S^{1}$-spaces are Kähler. In this article we consider a type of Hamiltonian $S^{1}$-action $\Phi:S^{1}\to Ham(M,\omega)$ on a symplectic 4-manifold $(M,\omega)$ with an extra property called the Maslov condition, which requests that along any nondegenerate orbit $C$, any $\Phi$-invariant Lagrangian subbundle in the symplectic normal bundle of $C$, treated as a loop of Lagrangian planes along $C$, has zero Maslov number. We call such a group action a special Hamiltonian $S^{1}$-action (SHam1-action) and the triple $(M,\omega,\Phi)$ a special Hamiltonian $S^{1}$-space (SHam1-space). See Definition 2.1 in §2 for the precise definition of SHam1-action. To ensure the Maslov condition is permitted we assume that $(M,\omega)$ is connected with the first Chern class $c_{1}(M,\omega)=0$, and that the SHam1-action is effective on $M$.
It turns out that the Maslov condition imposes restrictions on the topology of $M$ as well of the $S^{1}$-action.
@@
Let $(M,\omega,\Phi)$ be a connected simple SHam1-space and $\Phi$ acts effectively on $M$ with moment map $h$, then
@@
    the fixed point set $Fix(\Phi)$ is a discrete set of points (Proposition 3.4).
@@
    Moreover, if $(M,\omega=d\lambda)$ is exact then $Fix(\Phi)\subset h^{-1}(0)$ up to addition of a constant to $h$ (Proposition 3.5).
Since a SHam1-space $(M,\omega,\Phi)$ is not compact, if we remove a codimension $\geq 1$ $\Phi$-invariant subset, e.g., any number of nonconstant $\Phi$-orbits, from $M$, then $\Phi$ restricts to a SHam1-action on the remaining symplectic manifold $(M^{\prime},\omega)$, so $(M^{\prime},\omega,\Phi)$ by itself is also a SHam1-space, but the topology of $M^{\prime}$ can be arbitrarily complicated. To avoid extra topological complexity of $(M,\omega,\Phi)$ beyond the presence of a SHam1-action, we assume that $(M,\omega=d\alpha)$ is simple, which is defined as follows:
Definition 1.2.
Let $(M,\omega=d\lambda,\Phi)$ be a connected open exact SHam1-space, and let $h$ denote the moment map associated to $\Phi$. $(M,\omega=d\lambda,\Phi)$ is simple if
  1. (i).
    $M$ is a complete manifold with respect to some Riemannian metric, and is homotopic to a finite handlebody, or equivalently, a finite CW-complex,
  2. (ii).
    the gradient vector field $\nabla h$ of the moment map $h$ is complete with respect to some $\omega$-compatible $\Phi$-invariant Riemannian metric on $M$,
  3. (iii).
    level sets $h^{-1}(c)$ are connected for all $c\in\mathbb{R}$, and all reduced spaces $h^{-1}(c)/\Phi$ are homeomorphic to a noncompact surface without punctures.
In this paper we consider the classification problem of all (connected) simple exact SHam1-spaces $(M,\omega=d\lambda,\Phi,h)$ with finite fixed point set $Fix(\Phi)$. All SHam1-sapces considered in this manuscript are assumed to be simple.
For the 1-connected case we obtain the following topological classification results.
@@
Let $(M,\omega,\Phi)$ be a connected, 1-connected simple exact SHam1-space, and the $\Phi$-action is semi-free with $n+1$ fixed points for some integer $n\geq 0$. If $n>0$ then $(M,\omega)$ is symplectically the linear plumbing of $n$ cotangent bundles of spheres $S_{i}\subset M$, i=1,,ni=1,...,n, where the union $\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of $\Phi$-invariant Lagrangian spheres. If $n=0$ then $(M,\omega)$ is diffeomorphic to the standard symplectic 4-space $(\mathbb{R}^{4},\omega_{0}=\sum_{i=1^{2}}(dx_{i}\wedge dy_{i})$.
@@
Let $\omega_{n}$ denote the standard symplectic structure on $\mathbb{C}^{3}$ restricted to $\mathcal{W}_{n}$. It is known that for $n>0$ $\mathcal{W}_{n}$ is topologically the plumbing of $n$ copies of cotangent bundles $T^{*}S^{2}$ (of Lagrangian spheres) of type $A_{n}$ [16], and $\mathcal{W}_{0}$ is symplectomorphic to the standard symplectic 4-space.
@@
Let $\tilde{\Theta}$ denote the $S^{1}$-action on $\mathbb{C}^{3}$ defined by
@@
The action of $\tilde{\Theta}$ preserves $\mathcal{W}_{n}$. We denote by $\Theta$ the induced action on $\mathcal{W}_{n}$. $\Theta$ acts on $(\mathcal{W}_{n},\omega_{n})$ as a SHam1-action with $n+1$ Fixed points, and $\textsf{h}:=x_{2}y_{1}-x_{1}y_{2}$ is the moment map with $Fix(\Theta)\subset\textsf{h}^{-1}(0)$.
Remark 1.5.
Note that $(\mathcal{W}_{n},\omega_{n}=d\alpha,\Theta^{-1})$ is also a SHam1-space, where $\Theta^{-1}$ is the inverse action of that of $\Theta$. More generally, if $(M,\omega,\Phi)$ is a SHam1-space, then so is $(M,\omega,\Phi^{-1})$. We will show in Proposition 3.5 that for any connected, exact, simple effective SHam1-space $(M,\omega,\Phi)$, the weight of the $\Phi$-action at its fixed points are either all equal to 1, or all equal to -1, relative to the Darboux charts at each of the fixed points.
Definition 1.6.
Two SHam1-spaces $(M_{i},\omega_{i},\Phi_{i})$, i=1,2=1,2 are symplectically equivariant if there exists a diffeomorphism $f:M_{1}\to M_{2}$ such that $f^{*}\omega_{2}=\omega_{1}$ and $^{*}\Phi_{2}=\Phi_{1}. ff$ is called an isomorphism between the two SHam-1 spaces as in [8].
Proposition 1.7.
Let $\Theta^{-1}$ denote the inverse SHam1-action of $\Theta$ on $(\mathcal{W}_{n},\omega_{n})$. Then there exists a Hamiltonian diffeomorphism $f$ of $(\mathcal{W}_{n},\omega_{n})$, such that $f^{-1}\circ\Theta^{-1}\circ f=\Theta$, i.e., actions of $\Theta$ and $\Theta^{-1}$ on $(\mathcal{W}_{n},\omega_{n})$ are equivariant up to a conjugation by a Hamiltonian diffeomorphism on $(\mathcal{W}_{n},\omega_{n})$. In particular, we can take
$f(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3}):=f(-x_{1},-y_{1},x_{2},y_{2},x_{3},y_{3}),$
then $f=f^{-1}$ and $f$ preserves the Lefschetz fibers of $\Theta$ (hence of $\Theta^{-1})$. See Remark 2.3 and Remark 5.1 for detail.
@@
A connected, 1-connected, simple SHam1-space $(M,\omega=d\lambda,\Phi)$ with $n+1$ fixed points is symplectomorphic to the Stein surface $(\mathcal{W}_{n},\omega_{n})$. In particular, the symplectic topology of $(M,\omega)$ is determined by the cardinality of the fixed point set $Fix(\Phi)$.
Moreover, together with the following theorem by Wu [16], any $A_{n}$-string of Lagrangian spheres in $(M,\omega)$ is Hamiltonian isotopic to any given $A_{n}$-string of Lagrangian spheres associated to the $\Theta$-action as in Theorem 1.8, up to a composition of Lagrangian Dehn twists along the latter:
@@
Any compactly supported symplectomorphism of $\mathcal{W}_{n}$ is Hamiltonian isotopic to a composition of Dehn twists along the standard spheres. In particular, $\pi_{0}(\text{Symp}_{c}(\mathcal{W}_{n}))=Br_{n+1}$. As a result, exact Lagrangians in $A_{n}$-surface singularities are isotopic to the zero section of a plumbed copy of $T^{*}S^{2}$, up to a composition of Lagrangian Dehn twists along the standard spheres.
We have the following result concerning the uniqueness of special Hamiltonian $S^{1}$-actions on $\mathcal{W}_{n}$.
Theorem 1.10.
(Linearity and symplectic equivariance of SHam1-actions) Let $\Phi$ be any semi-free SHam1-action on $(\mathcal{W}_{n},\omega_{n})$ with moment map $h$ with $n+1$ fixed points, $n\in\mathbb{Z}$, $n\geq 0$, and let $S=S_{\phi}=\cup_{i=1}^{n}S_{i}\subset\mathcal{W}_{n}$ be the associated $A_{n}$-string of Lagrangian spheres. Also let $\textsf{S}=\textsf{S}_{\Theta}$ denote the $A_{n}$-string of Lagrangian spheres in $W_{n}$ associated to $\Theta$.
  1. (i).
    There is a $\Phi$-invariant open neighborhood $U\subset\mathcal{W}_{n}$ of S (resp. the fixed point $(0,0,1)\in\mathcal{W}_{0}$ if $n=0$) and a $\Phi$-equivariant $C^{\infty}$-diffeomorphism $\EuFrak{e}_{\textsf{S}}:U\to V=\EuFrak{e}_{\textsf{S}}(U)\subset\mathcal{W}_{n}$ on which $\Phi=\EuFrak{e}_{\textsf{S}}\circ\Theta\circ\EuFrak{e}_{\textsf{S}}^{-1}$, $\omega^{\prime}:=\EuFrak{e}_{\textsf{S}}^{*}\omega_{n}$ is symplectic and $\Phi$-invariant, the $\Phi$-action on $U$ is linear with respect to the symplectic form $\omega^{\prime}$. Then by Definition 1.6 $(U,\omega^{\prime},\Phi,h)$ and $(V,\omega_{n},\Theta,\textsf{h})$ are symplectically equivariant via the symplectomorphism $\EuFrak{e}_{\textsf{S}}:(U,\omega^{\prime})\to(V,\omega_{n})$.
  2. (ii).
    If $n=0$, any semi-free SHam1-action $\Phi$ on $(\mathbb{C}^{2},\omega)$ with $Fix_{\Phi}=\{0\}$ is linear up to conjugation and symplectically equivariant to the standard linear action of $\Theta$ on open 4-ball $B_{r}(0)=\{q\in\mathbb{C}^{2}\mid|q|^{2}<r\}$ centered at 0 with any prescribed radius $\sqrt{r}$.
  3. (iii).
    If $n=1$, then $\Phi$ is symplectically equivariant to $\Theta$ on any compactly supported $\Phi$-invariant open domain containing the $A_{1}$-sphere $S\cong S^{2}$ associated to the $\Theta$-action on $\mathcal{W}_{1}$.
For the case that $(M,\omega=d\lambda,\Phi,h)$ is not 1-connected we assume that $Fix(\Phi)$ is not empty (otherwise the $\Phi$-action is free and the Maslov condition is vacuous). Then the orbit space $h^{-1}(0)/\Phi$ is topologically an oriented noncontractible surface of genus $g\geq 0$ with $b\geq 1$ boundary components, $g+b\geq 2$, and $k:=|Fix(\Phi)|=n+1\geq 1$ marked points.
Theorem 1.11.
Let $(M,\omega,\Phi)$ be a connected exact SHam1-space with $c_{1}(M,\omega)=0$ and $n+1$ fixed points, $n\geq 0$. Assume also that $M$ is not 1-connected. Then $M$ admits the structure of a Stein surface, which is topologically obtained by attaching (2g+b1)(2g+b-1)-pairs of 1- and Lagrangian 2-handlebodies to the boundary of a Stein domain $W_{n}\subset\mathcal{W}_{n}$ diffeomorphic to $\mathcal{W}_{n}$. The homology groups of $M$ are
@@
In particular, $H_{2}(M,\mathbb{Z})$ is generated by $n$ Lagrangian spheres and 2g+b12g+b-1 monotone Lagrangian tori.
Remark 1.12 (Relation with complexity one Hamiltonian space).
The 4-dimensional exact SHam1-spaces $(M,\omega=d\lambda,\Phi)$ that we consider here are also centered complexity one spaces as defined in [9] (see also [10, 11, 12]) except that the moment maps of these spaces are not proper unless restricted to a $\Phi$-invariant domain $D$ of $M$, so that for any compact interval $a,b]\subset\mathbb{R}$ the preimage $h^{-1}([a,b])$ of the moment map is compact.
@@
The rest of this paper is organized as follows: In §2 we give the definition of a SHam1-action on a symplectic 4-manifold $(M,\omega)$. As an example we show that an $S^{1}$-subgroup of $SU(2)$ acts on the standard symplectic 4-space $(\mathbb{R}^{4},\omega_{0})$ special Hamiltonianly. Proof of Theorem 1.1 is given in §3 where we also establish various symplectic topological properties of SHam1-spaces. Theorem 1.3 is proved in §4. Proposition 1.4 and Theorem 1.8 are verified in §5 where the $\Theta$-action is analyzed and an $A_{n}$-string of Lagrangian spheres is given. This is followed by the proof of Theorem 1.10. Theorem 1.11 is proved in §6.